OpenStudy (anonymous):

How would you simplify this expression: 32^3/5?

OpenStudy (anonymous):

I think I've found the solution. Using this website as a guide http://www.mathwarehouse.com/algebra/exponents/fraction-exponents/formula-examples-simplify-exponents-with-fractions.php If I do:\[32^{\frac{ 1 }{ 5 }} \times 32^{\frac{ 1 }{ 5 }} \times 32^{\frac{ 1 }{ 5 }} = 2 \times 2 \times 2 = 8\]